Researchers measured the data speeds for a particular smartphone carrier at 50 airports. The highest speed measured was 71.2 Mbps. The complete list of 50 data speeds has a mean of x overbarequals18.77 Mbps and a standard deviation of sequals17.46 Mbps. a. What is the difference between​ carrier's highest data speed and the mean of all 50 data​ speeds? b. How many standard deviations is that​ [the difference found in part​ (a)]? c. Convert the​ carrier's highest data speed to a z score. d. If we consider data speeds that convert to z scores between minus2 and 2 to be neither significantly low nor significantly​ high, is the​ carrier's highest data speed​ significant? a. The difference is nothing Mbps.
Answer #1

a) Difference = 71.2 - 18.77 = 52.43

b) Number of standard deviations = 52.43/17.46 = 3.00

c) z = 3.00

d) Significantly High